\begin{thebibliography}{10}

\bibitem{ATV1}
M.~Artin, J.~Tate, and M.~Van~den Bergh.
\newblock Some algebras associated to automorphisms of elliptic curves.
\newblock {\em The {G}rothendieck {F}estschrift, {V}ol. {I}}, 86:33--85, 1990.

\bibitem{ATV2}
M.~Artin, J.~Tate, and M.~Van~den Bergh.
\newblock Modules over regular algebras of dimension {$3$}.
\newblock {\em Invent. Math}, 106(2):335--388, 1991.

\bibitem{AS}
M.~Artin, J.~Tate, and M.~Van~den Bergh.
\newblock Modules over regular algebras of dimension {$3$}.
\newblock {\em Invent. Math.}, 106(2):335--388, 1991.

\bibitem{AVdB1990}
M.~Artin and M.~Van~den Bergh.
\newblock Twisted homogeneous coordinate rings.
\newblock {\em J. Algebra}, 133(2):249--271, 1990.

\bibitem{MSRIBook}
G.~Bellamy, D.~Rogalski, T.~Schedler, J.~T. Stafford, and M.~Wemyss.
\newblock {\em Noncommutative algebraic geometry}, volume~64 of {\em
  Mathematical Sciences Research Institute Publications}.
\newblock Cambridge University Press, New York, 2016.
\newblock Lecture notes based on courses given at the Summer Graduate School at
  the Mathematical Sciences Research Institute (MSRI) held in Berkeley, CA,
  June 2012.

\bibitem{EH}
D.~Eisenbud and J.~Harris.
\newblock {\em The geometry of schemes}, volume 197 of {\em Graduate Texts in
  Mathematics}.
\newblock Springer-Verlag, New York, 2000.

\bibitem{Hart}
R.~Hartshorne.
\newblock {\em Algebraic geometry.}
\newblock Springer-Verlag, New York-Heidelberg, 1977.

\bibitem{lBSV}
L.~Le~Bruyn, S.~P. Smith, and M.~Van~den Bergh.
\newblock Central extensions of three-dimensional {A}rtin-{S}chelter regular
  algebras.
\newblock {\em Math. Z.}, 222(2):171--212, 1996.

\bibitem{Le}
T.~Levasseur.
\newblock Some properties of noncommutative regular graded rings.
\newblock {\em Glasgow Math. J.}, 34(3):277--300, 1992.

\bibitem{SV99}
B.~Shelton and M.~Vancliff.
\newblock Embedding a quantum rank three quadric in a quantum {${\bf P}^3$}.
\newblock {\em Comm. Algebra}, 27(6):2877--2904, 1999.

\bibitem{Si}
S.~J. Sierra.
\newblock {$G$}-algebras, twistings, and equivalences of graded categories.
\newblock {\em Algebr. Represent. Theory}, 14(2):377--390, 2011.

\bibitem{HTDiss}
H.~V. Tran.
\newblock Twisting systems and some quantum ${P}^3$s with point scheme a rank-2
  quadric.
\newblock {\em Dissertation}, 2022.

\bibitem{HV}
H.V. Tran and M.~Vancliff.
\newblock Twisting systems and some quantum ${P}^3$s with point scheme a rank-2
  quadric.
\newblock In K.A.~Brown et~al, editor, {\em Contemporary Math}. Amer. Math.
  Soc., to appear.

\bibitem{VVr}
M.~Vancliff and K.~Van~Rompay.
\newblock Embedding a quantum nonsingular quadric in a quantum {${\bf P}^3$}.
\newblock {\em J. Algebra}, 195(1):93--129, 1997.

\bibitem{Zhang96}
J.~J. Zhang.
\newblock Twisted graded algebras and equivalences of graded categories.
\newblock {\em Proceedings of the London Mathematical Society. Third Series},
  72(2):281--311, 1996.

\end{thebibliography}

\documentclass[12pt,reqno]{amsart}
\usepackage{amsfonts,amsmath,amssymb,amsthm,mathrsfs,textcomp,enumerate,comment,cleveref,breakurl}

\usepackage{pdfpages}
\usepackage{lscape}
\usepackage{multicol}

\DeclareMathOperator{\Aut}{Aut}
\DeclareMathOperator{\Gr}{Gr}
\DeclareMathOperator{\QGr}{QGr}
 
\DeclareMathOperator{\gldim}{gldim}

\DeclareMathOperator{\id}{id}

\renewcommand{\k}{{\Bbbk}}
\renewcommand{\P}{\mathbb{P}}

\newcommand{\kk}{\Bbbk}

\newcommand{\V}{\mathcal{V}}

\newtheorem{thm}{Theorem}[section]
\newtheorem{lemma}[thm]{Lemma}
\newtheorem{cor}[thm]{Corollary}
\newtheorem{prop}[thm]{Proposition}

\theoremstyle{definition}

\newtheorem{notn}[thm]{Notation}
\newtheorem{ex}[thm]{Example}
\newtheorem{rmk}[thm]{Remark}

\newtheorem{example}[thm]{Example}

\newtheorem{Pf}{Proof$\!\!$}

\DeclareMathSymbol{\twoheadrightarrow}  {\mathrel}{AMSa}{"10}
\newcounter{letter}
\renewcommand{\theletter}{\rom{(}\alph{letter}\rom{)}}

\newcounter{rnum}
\renewcommand{\thernum}{\rom{(}\roman{rnum}\rom{)}}

\usepackage{xcolor}
\usepackage[normalem]{ulem}

\begin{document}
%\baselineskip21pt

%\hspace*{\fill}\today

\title[Rank 2 Quadric]%
{Regular Algebras of Dimension Four Associated to Coordinate Rings of Rank-Two Quadrics}
\subjclass[2020]{14A22,16S37,16S38}
%\subjclass[2010]{16S37, 16S38}%
%14A22 Noncommutative algebraic geometry
%16S37 Quadratic and Koszul algebras
%16S38 Rings arising from non-commutative algebraic geometry
%17B35 Universal enveloping algebras
%17B37 Quantum enveloping algebras
%17B75 Color Lie algberas

\keywords{Artin-Schelter regular algebra, point scheme, coordinate rings, rank-two quadric%
%\newline
%\indent This work was supported in part by NSF grants DMS-0900239 and
%DMS-1302050
}
%\rule[-5mm]{0cm}{0cm}}%

\maketitle

\vspace*{0.1in}

\renewcommand{\thefootnote}{\fnsymbol{footnote}}

\begin{center}
\textsc{Richard G.\ Chandler}\\
richard.chandler@untdallas.edu,\\
 
 \bigskip
 
\textsc{Hung Tran}\\
hvtran01@gmail.com,\\

\bigskip

\textsc{Padmini Veerapen}\\
pveerapen@tntech.edu,\\

\medskip
and
\medskip

\textsc{Xingting Wang}\\
xingtingwang@math.lsu.edu

\end{center}

\setcounter{page}{1}
\thispagestyle{empty}

\bigskip
\bigskip

\begin{abstract}

In this paper, we classify connected graded quadratic Artin-Schelter regular (AS-regular, henceforth) algebras of global dimension four that have a Hilbert series the same as that of the polynomial ring on four generators and that map onto a twisted homogeneous coordinate ring of a rank-two quadric. A twisted homogeneous coordinate ring is a construction that was defined by Artin, Tate, and Van den Bergh in \cite{ATV1, ATV2, AVdB1990} in the context of the classification of AS-regular algebras of global dimension three. In \cite{SV99,VVr}, the authors classified AS-regular algebras of global dimension four that map onto the twisted homogeneous coordinate ring of a rank-three and a rank-four quadric, respectively. We expand on their work to include the case of coordinate rings of a rank-two quadric.
\end{abstract}

\baselineskip18pt

%\newpage
\bigskip
\bigskip

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
% INTRO
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\section*{Introduction}
It is widely accepted that Artin-Schelter regular (AS-regular, henceforth) algebras can be viewed as noncommutative analogues of polynomial rings.  These algebras were defined by  Artin and Schelter in 1987 \cite{AS}; Artin, Tate and Van den Bergh \cite{ATV1, ATV2} classified those of global dimension three using techniques that became known as noncommutative projective algebraic geometry. In this paper, our overarching goal is to provide examples that would contribute to the classification AS-regular algebras of dimension four by utilizing these geometric techniques.

In \cite{AVdB1990}, Artin and Van den Bergh elaborate on the construction of a ``noncommutative" version of the commutative homogeneous coordinate ring, known as a twisted homogeneous coordinate ring. As defined in \cite[Definition p. 249]{AVdB1990}, a twisted homogeneous coordinate ring, $S$, is given by three pieces of data, a projective scheme $Q$ defined over a field $\Bbbk$, an automorphism $\sigma$ of $Q$, and an invertible sheaf on $Q$, that is, we may write $S$ as $S(Q, \sigma, \mathcal{L})$. In \cite[Definition 4.3.1]{MSRIBook}, Rogalski highlights the fact that if the invertible sheaf $\mathcal{L}$ is $\sigma$-ample, then such rings are noetherian and finitely generated by \cite[Theorem 1.4]{AVdB1990}.

Shelton, Vancliff, and Van Rompay further examine the twisted homogeneous ring construction in \cite{SV99, VVr} by classifying the quadratic noetherian AS-regular algebras of dimension four that map onto the twisted homogeneous coordinate ring of a rank-three and a rank-four quadric in $\mathbb{P}^3$.  The algebras are classified by considering whether their point schemes $P$ either strictly contains a rank-two quadric, that is, $Q \subset P,$ or are exactly a rank-two quadric, that is, $P = Q$.  In the case of the quadric being strictly contained in the point scheme, it was shown that the algebra is completely determined by the geometric data and, thus, the defining relations can be directly recovered. In this paper, we extend their work to twisted homogeneous coordinate rings of a rank-2 quadric, $Q \subset \mathbb{P}^3$.

In Section \ref{sec1}, Theorem \ref{OmegaNormal}, we prove a result critical to the rest of the paper. Assuming that $R$ is a quadratic domain with $\Omega \in R_2$ and such that $A/{\langle \Omega \rangle} = S$, we show that $\Omega$ is normal. In Section \ref{sec2}, we identify regular algebras of global dimension four whose point scheme strictly contains a rank-two quadric.  We begin by finding such algebras whose automorphism on the quadric is the identity and then twist these algebras by a non-identity automorphism in $\Aut(\P^3)$; that is, we twist them via a Zhang twist. In Section \ref{sec3}, we show, in Proposition \ref{prop1}, that when the point scheme of $R$ is exactly a rank-two quadric, then the quadratic noetherian AS-regular algebras of dimension four are, in fact, central extensions of  AS-regular algebras of dimension three. 

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
% Preliminaries
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\section{Preliminaries}\label{sec1}

\begin{notn}
Let $Q \in \P^3$ be a rank-two quadric and fix an automorphism $\tau$ of $Q$. By $S(Q)$, we mean the (commutative) homogeneous coordinate ring of $Q$. By $S$, we mean the twisted homogeneous coordinate ring corresponding to a rank-2 quadric $Q$, an automorphism $\tau$ of $Q$ and a $\tau$--ample invertible sheaf $\mathcal{L}$, that is, $S = S(Q, \tau, \mathcal{L})$. The set $\mathcal{R}(Q, \tau)$ consists of all graded algebras $R$ that are quadratic connected Artin-Schelter regular of global dimension four with Hilbert series $1/(1-t)^{-4}$ and such that there is a graded degree zero homomorphism mapping $R$ onto $S$. We denote $R \in \mathcal{R}(Q, \tau)$ by $T(V)/I$ where $V$ is a four-dimensional vector space, $T(V)$ is the tensor algebra on $V$ and $I$ is a finitely generated graded ideal. The graph of $\tau$ on $Q$ will be denoted by $\Gamma_{\tau}(Q)$.
\end{notn}

%\begin{lemma}\cite[Lemma 1.3]{SV99}\label{6QuadRelns}
In \cite[Lemma 1.3]{SV99}, if $A$ is Artin-Schelter regular of global dimension four with four linearly independent generators, then $A$ is Koszul and has Hilbert series $H_A(t) = (1-t)^{-4}$. In particular, $A$ is quadratic with six defining relations and has Gelfand-Kirillov dimension four. Henceforth, we write $R$ as $R = T(V)/W$ where $W$ is a six-dimensional vector space spanned by the defining relations of $R$. 

\begin{rmk}\label{omega}
\begin{enumerate}
\item Using an argument similar to \cite[Lemma 1.7]{SV99}, we note that the Hilbert series of a twisted homogeneous coordinate ring $S$ of any quadric $Q' \in \P^3$ is $(1+t)(1-t)^{-3}$. This follows from the fact that we may view $S$ as a quotient of $S(\P^3)$ by the ideal generated by $Q'$. As a result, the Hilbert series of $S$ is that of $Q'$ subtracted from that of $S(\P^3)$ which is given by $(1-t)^{-4} - t^2(1-t)^{-4}$. In general, we will use $S($--$)$ to mean the (commutative) homogeneous coordinate ring corresponding to that scheme.
\item By hypothesis, $R$ has Hilbert series $(1+t)^{-4}$ and maps onto $S$ with Hilbert series $(1+t)(1-t)^{-3}$. Similar to \cite[before Definition 1.3]{VVr}, this implies that there exists $\Omega \in R_2$ such that $R/{\langle \Omega \rangle} \cong S$. Now, we know that there are no nonzero elements of $S_2$ that vanish on the graph of $\tau$, as $S$ is a domain, so any nonzero element of $R_2$ that vanishes on $\tau$ must be a scalar multiple of $\Omega$. Thus, for any $u, v \in R_1$, $u^{\tau}v - v^{\tau}u \in R_2$ and is a scalar multiple of $\Omega$.    
\end{enumerate}
\end{rmk}

\begin{example}\label{omegaNotNormal}
Let $A$ be the $\k$-algebra given by the polynomial ring $\k[x_1,x_2,x_3,x_4]$ subject to the following defining relations $$\begin{matrix} x_1x_2-x_2x_1, & x_1x_3-x_3x_1, & x_1x_4-x_4x_1, & x_2x_4-x_4x_2, & x_3x_4-x_4x_3, & x_1x_2. \end{matrix}$$
This algebra is not noetherian since the following chain of right ideals does not stablize: $$x_2A \subset x_2A + x_3x_2A \subset x_2A+x_3x_2A+x_3^2x_2A\subset \cdots \subset \sum_{i=0}^n x_3^i x_2 A \subset \cdots$$
However, $A/\langle \Omega \rangle$, where $\Omega=x_2x_3-x_3x_2 \in A$, is the coordinate ring of the rank-two quadric $\V(x_1x_2)$, which is noetherian.  Therefore, \cite[Lemma 8.2]{ATV1} implies that $\Omega$ is not normal in $A$.
\end{example}

\begin{rmk}\label{zhangtwist}
\begin{enumerate}
\item By Example \ref{omegaNotNormal}, we see that, in general, for a quadratic element $\Omega \in R_2$, $R/{\langle \Omega \rangle} \cong S$ does not ensure that $\Omega$ is normal. This is in agreement with \cite[Example 1.6]{SV99} when the quadric is rank-three but differs from \cite[Lemma 1.5]{VVr} when the quadric is rank-four.
\item 
By \cite[Theorems 3.1 \& 3.5]{Zhang96}, when a connected graded algebra $A$ is twisted by an automorphism or by a twisting system of $A$ denoted by $\phi$, then the category of graded modules of $A$, $\Gr A$, and that of graded modules over $A^{\phi}$, $\Gr A^{\phi}$, are equivalent; that is, the category of graded modules of $A$ are invariant under twists by automorphisms and twisting systems. Additionally, by \cite[Theorem 1.4]{Zhang96}, since $\QGr A$ is equivalent to $\QGr A^{\phi}$, it follows that the point schemes of $A$ and $A^{\phi}$ are equivalent, where $\QGr A$, is the quotient category $\Gr A/\mbox{tors } A$ and tors $A$ is the full subcategory of $\Gr A$ consisting of modules which are finite dimensional. The global dimension and Hilbert series of $A$ and $A^{\phi}$ also remain invariant. We refer to the algebras $A$ and $A^{\phi}$ as \textit{Zhang twists} of each other as suggested by other authors (see e.g., \cite{Si}).
\end{enumerate}
\end{rmk}

\begin{prop} \label{OmegaNormal} If $A$ is a quadratic domain  with $\Omega \in A_2$ such that $A/\langle \Omega \rangle = S$, then $\Omega$ is normal in $A$. 
\begin{proof}
In this proof, we use techniques similar to \cite[Lemma 1.7]{SV99} with significant modifications. By Remark \ref{omega}, we may assume that $\tau = \id$, that is, $S=S(Q)$,  and $uv-vu\in \k \Omega$ for all $u,v \in A_1$.
Let $Q=\V(X_1X_2)$, where $X_1, X_2 \in A_1$ are linearly independent, and $X_3 \in A_1$ such that $X_3 \notin \Bbbk^{\times} X_1 \oplus \Bbbk^{\times} X_2$. It follows that $X_iX_j-X_jX_i=\alpha_{ij}\Omega$ for $\alpha_{ij} \in \k$, $i,j \in \{1,2,3\}$.  Define $y=\alpha_{23}X_1 - \alpha_{13}X_2+\alpha_{12}X_3$.  It can be verified that $y$ commutes in $A$ with all elements of the form $\k X_1+\k X_2+\k X_3$. Note also that $X_1X_2\in \k^\times \Omega$, as it is an element of $A_2$ that vanishes on the graph of $Q$ and is nonzero since $A$ is a domain; thus, $\Omega\in\k^\times X_1X_2$. Observe that the plane $\V(y)\subset \P^3$ meets $Q$ either in two lines, counted with multiplicity, or in a plane.

If $\V(y)$ meets $Q$ in two lines, counted with multiplicity, then we may present $A$ using generators $x_1=X_1$, $x_2=X_2$, and $x_3=y$ and so, $x_1x_3=x_3x_1$ and $x_2x_3=x_3x_2$. Moreover, if $\alpha_{12} = 0$, that is, if $x_1x_2=x_2x_1$, then $\Omega$ commutes with $x_1$, $x_2$ and $x_3$.  Additionally, if $x_4 \in A_1 \setminus (\k x_1 + \k x_2 + \k x_3)$, then
\begin{align*} x_4 \Omega
&= x_4(\beta x_1 x_2), \qquad \mbox{ where } \Omega = \beta x_1x_2 \\
&= \beta(x_1x_4 + \gamma \Omega) x_2 \\
%&= \beta (x_1x_4x_2 + \gamma \Omega x_2) \\ 
&= \beta (x_1(x_2 x_4 + \delta \Omega)+\gamma \Omega x_2) \\
%&= \beta(x_1x_2x_4 + \delta x_1 \Omega + \gamma \Omega x_2)\\
%&= \Omega x_4 + \beta \delta \Omega x_1 + \beta \gamma \Omega x_2\\ 
&= \Omega(x_4+\beta \delta x_1 + \beta \gamma x_2),  
\end{align*}
where $\gamma,\delta \in \k$, that it, $\Omega$ is normal in $A$. Otherwise, $\alpha_{12} \ne 0$, that is, $x_1x_2\neq x_2x_1$, and we show there is $x_4 \in A_1 \setminus (\k x_1 + \k x_2 + \k x_3$) such that $x_1x_4 = x_4 x_1$ and $x_2 x_4 = x_4 x_2$.  Choose $w \in A_1 \setminus (\k x_1 + \k x_2 + \k x_3$).  If $w$ commutes with $x_1$, define $z=w$; otherwise $x_1w-wx_1=\beta \Omega$, for $\beta \in \k^\times$.  In this case, define $z=\alpha_{12}w-\beta x_2$; in either case, $z$ commutes with $x_1$.  If, in addition, $z$ commutes with $x_2$, define $x_4=z$.  Otherwise, $x_2z-zx_2=\gamma \Omega$, for some $\gamma \in \k^\times$; define $x_4=\alpha_{12}z + \gamma x_1$. This $x_4$ commutes with $x_2$ (and with $x_1)$, so we may always find an $x_4$ as described above.  It follows that $\Omega$ is normal in $A$.

If $\V(y)$ meets $Q$ in a plane, then we may assume without loss of generality that $y = \alpha_{23} X_1$. That is, $\alpha_{13} = 0 = \alpha_{12}$ and the defining relations of $A$ include $X_1X_2=X_2X_1$, $X_1X_3=X_3X_1$, and $X_2X_3-X_3X_2=\alpha_{23}\Omega$.  Choose $w \in A_1 \setminus (\k X_1 + \k X_2 + \k X_3$).  If $w$ commutes with $X_3$, define $z=w$; otherwise $X_3w-wX_3=\beta \Omega$, for $\beta \in \k^\times$.  In this case, define $z=\alpha_{23}w+\beta X_2$; in either case, $z$ commutes with $X_3$. If, in addition, $z$ commutes with $X_2$, define $X_4=z$. Otherwise, $X_2z-zX_2=\gamma \Omega$, for some $\gamma \in \k^\times$; define $X_4=\alpha_{23}z - \gamma X_3$. As chosen, $X_4$ commutes with $X_2$ (and with $X_3)$, so we may always find an $X_4$ as described above.  It follows that $\Omega$ is normal in $A$. 
\end{proof}
\end{prop}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
% END  PRELIMINARIES
% START  SECTION  1
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\bigskip
%\newpage

\section{Regular Algebras whose Point Scheme is not the Union of Two Planes}\label{sec2}

\begin{rmk}\label{HypsOnR}
Since $R$ is AS-regular with Hilbert series $H_R(t) = (1-t)^{-4}$, by \cite[Definition in \S 5.8, Corollary in \S 6.2]{Le}, we may assume $A$ is a quadratic noetherian algebra satisfying
\end{rmk}
\begin{enumerate}
\item $H_A(t) = (1-t)^{-4}$;
\item $A$ is Auslander-regular of global dimension four;
\item $A$ satisfies the Cohen-Macaulay property.
\end{enumerate}

\begin{lemma}
Let $x_i$ and $x_j$ be two degree-one elements in $R$. If $x_i$ and $x_j$ are both normal in $R$, then $x_ix_j = \alpha_{ij} x_jx_i$, for some $\alpha_{ij} \in \k^\times$.
\end{lemma}
\begin{proof}
Since $x_1$ and $x_2$ are two normal elements in the quadratic algebra $R$, it follows that $x_1x_2 \in R_1 x_1 \cap x_2 R_1$ and $x_2x_1 \in R_1 x_2 \cap x_1 R_1$; hence, $x_1x_2 \in \k^{\times} x_2 x_1$ and $x_2x_1 \in \k^{\times} x_1 x_2$ as $R$ is a domain.    
\end{proof}

\begin{lemma}\cite[Lemma 2.2 \& Proposition 2.4]{VVr}\label{normalElt}
If the point scheme $P$ is not the quadric $Q$, then $W'$ is a five-dimensional subspace of $W$. Moreover, there exists a normal element in $R_1$.
\end{lemma}

\begin{prop}\cite[Lemma 2.5]{VVr}
\label{normalautom}
If $\omega \in R_1$ is normal and $\phi \in \Aut(R)$ is defined by $\omega x = x^{\phi}\omega$ for all $x \in R$, then $\phi \in \kk^{\times} \tau$.     
\end{prop}

\begin{prop}\cite[Theorem 2.6]{VVr}\label{centralElts}
If the point scheme $P$ of $R$ is not the quadric $Q$, then $\tau \in \Aut(R)$. In this case, $R$ is a Zhang twist of an AS-regular algebra $R'$ that maps onto the (commutative) homogeneous coordinate ring $S(\P^3)$. In particular, $R'$ contains two linearly independent central elements of degree one.
\end{prop}
\begin{proof}
If $P \ne Q$, then $P$ contains a normal element by Lemma \ref{normalElt}, so by Proposition \ref{normalautom} $\tau \in \Aut(R)$. As a result, one may twist $R$ by $\tau^{-1}$ to obtain an AS-regular algebra $A'$ of $\gldim$ 4 (see, Remark \ref{zhangtwist}). Thus, the automorphism $\tau'$ of $Q$ corresponding to $R'$ is $\tau \circ \tau^{-1}$, which is the identity (cf. \cite[\S 1.3]{SV99}). That is, $R'$ maps onto $S(\P^3)$. 

Now, let $v_1, v_2 \in R_1$ be linearly independent. By Lemma \ref{normalElt}, the span $W'$ of defining relations of $R'$ contains a five-dimensional subspace generated by $\{v_i \otimes x - x \otimes v_i: x \in R_1', i = 1, 2\}$, which proves that $v_1$ and $v_2$, are central.   
\end{proof}

\begin{rmk}
One may characterize any scheme as being reduced or nonreduced, (see e.g., \cite[Example II.9]{EH} and \cite[Definition \S II.3]{Hart}). We denote the reduced scheme of $P$ by $P_{\mbox{red}}$.
\end{rmk}

\begin{prop}\cite[Proposition 2.7]{VVr} \label{quadline}
If the point scheme $P$ of $R$ is not the quadric $Q$, then either
\begin{enumerate}[(a)]
\item	$P=\P^3$
\item	$P=Q \cup L$ where $L$ is a line in $\P^3$ that meets $Q$ in two points (counted with multiplicity), or 
\item	the reduced scheme $P_{red}$ associated to $P$ is the quadric $Q$ but the scheme $P$ contains a double line $L$ of (all) multiple points on $P$, where $L$ corresponds to a line on $Q$ (we denote this situation by $P=Q\uplus L$).		
\end{enumerate}	
	
In cases (b) and (c) there is a regular normalizing sequence $\{v_1,v_2\} \subset R_1$ such that $L=\V(v_1,v_2)$.
\end{prop}	
%\begin{proof}
%The proof is equivalent to that of Proposition 2.7 in \cite{VVr}.
%\end{proof}

\begin{prop} \label{PLAlg}
Suppose that the point scheme $P$ of $R$ is not $Q$. If $\sigma|_Q = \tau = \id$, then there is a choice of generators $x_1, x_2, x_3, x_4$ for $R$ such that $Q=\V(x_1x_2)$ with the following six quadratic defining relations.
\begin{enumerate}[(a)]
\item $R=S(\P^3)=\k[x_1,x_2,x_3,x_4]$ is the polynomial ring on four variables and $(P,\sigma)=(\P^3,\id)$.
\item	$R=\k[x_1,x_2,x_3,x_4]$ with defining relations
$$\begin{matrix} x_3 x_1 = x_1 x_3, & x_4 x_2 = x_2 x_4, \\ x_3 x_2 = x_2 x_3, & x_4 x_1 = x_1 x_4, \\ x_3 x_4 = x_4 x_3, & x_2 x_1 = (\alpha - 1) x_1 x_2,  \end{matrix}$$
where $\alpha \in \k^\times \setminus\{1,2\}$.  In this case, $P=Q\cup L$ where $L=\V(x_3,x_4)$ (so $L$ intersects $Q$ at two distinct points) and $\sigma|_L(x_1,x_2,0,0)=\left( (\alpha-1)x_1, x_2,0,0 \right)$.
\item	$R=\k[x_1,x_2,x_3,x_4]$ with defining relations
$$\begin{matrix} x_3 x_1 = x_1 x_3, & x_2 x_1 = x_1 x_2, \\ x_3 x_2 = x_2 x_3, & x_2 x_4 - x_4 x_2 = x_1 x_2, \\ x_3 x_4 = x_4 x_3, & x_1 x_4 - x_4 x_1 = x_1 x_2.  \end{matrix}$$		
In this case, $P=Q\cup L$ where $L=\V(x_1-x_2,x_3)$ (so $L$ intersects $Q$ at one point with multiplicity two) and $\sigma|_L(x_1,x_1,0,x_4)=\left(x_1,x_1,0, x_1+x_4\right)$.
\item	$R=\k[x_1,x_2,x_3,x_4]$ with defining relations
$$\begin{matrix} x_1 x_2 = x_2 x_1, & x_2 x_3 = x_3 x_2, \\ x_1 x_3 = x_3 x_1, & x_2 x_4 = x_4 x_2, \\ x_1 x_4 = x_4 x_1, & x_3 x_4 - x_4 x_3 = x_1 x_2.  \end{matrix}$$
In this case, $P=Q\uplus L$ where $L=\V(x_1,x_2) \subset Q$ is the line of intersection of the planes of $Q$ and $\sigma \in \mathrm{Aut}(P)$ is uniquely determined by its action on the localized homogeneous coordinate rings of $P$, $S(P)[x_i^{-1}]$,  as follows: on $S(P)[x_1^{-1}]$ and $S(P)[x_2^{-1}]$ we have that $\sigma$ is the identity, but on $S(P)[x_3^{-1}]$ we have $$\sigma(x_1,x_2,x_3,x_4)=\left(x_1,x_2,x_3,x_4-x_3^{-1}x_1x_2\right),$$ and on $S(P)[x_4^{-1}]$ we have $$\sigma(x_1,x_2,x_3,x_4)=\left( x_1,x_2,x_3+x_4^{-1}x_1x_2,x_4 \right).$$
\item	$R=\k[x_1,x_2,x_3,x_4]$ with defining relations
$$\begin{matrix} x_1 x_2 = x_2 x_1, & x_2 x_3 = x_3 x_2, \\ x_1 x_3 = x_3 x_1, & x_3 x_4 = x_4 x_3, \\ x_1 x_4 = x_4 x_1, & x_2 x_4 - x_4 x_2 = x_2 x_1.  \end{matrix}$$
In this case, $P=Q\uplus L$ where $L=\V(x_1,x_3) \subset \V(x_1) \subset Q$ and $\sigma \in \mathrm{Aut}(P)$ is uniquely determined by its action on the localized homogeneous coordinate  rings of $P$, $S(P)[x_i^{-1}]$, as follows: on $S(P)[x_1^{-1}]$ and $S(P)[x_3^{-1}]$ we have that $\sigma$ is the identity but on $S(P)[x_2^{-1}]$ we have $$\sigma(x_1,x_2,x_3,x_4)=\left( x_1,x_2,x_3,x_1+x_4 \right),$$ and on $S(P)[x_4^{-1}]$ we have $$\sigma(x_1,x_2,x_3,x_4)=(x_1,x_2-x_4^{-1}x_2x_1,x_3,x_4).$$
\end{enumerate}
\end{prop}	
\begin{proof}
Our hypotheses imply that $R$ maps onto the (commutative) homogeneous coordinate ring, $S(Q)$.  As a result, if $x_1, x_2, x_3$, and $x_4$ are generators for $S(Q)$ such that $Q=\V(x_1x_2)$, then the defining relations of $R$ must be nonzero linear combinations of $x_1x_2$ and $x_ix_j-x_jx_i$ for all $i,j$, by \Cref{omega}.
\begin{enumerate}[(a)]
\item If $P=\P^3$, then $\sigma \in \mathrm{Aut}(\P^3)$ is linear. Thus, $\sigma = \tau =$ identity and $R=S(\P^3)$.
\item	If $L$ is a line in $\P^3$ that meets $Q$ in two distinct points, then we may choose coordinates $x_1, x_2, x_3, x_4$ such that $L=\V(x_3,x_4)$ and $Q=\V(x_1x_2)$.  By Proposition \ref{centralElts}, we may choose $x_3$ and $x_4$ to be central elements in $R$ which yields five defining relations for $R$.  We know that $x_1x_2-x_2x_1 \neq 0$ in $R$ as $P\neq \P^3$. Thus, $\alpha x_1 x_2 = x_2x_1-x_1x_2$ for some $\alpha \in \k^\times$.  This yields the sixth defining relation, $x_2 x_1 = (\alpha-1) x_1 x_2$. We note that $\alpha \ne 1$, as $R$ is a domain.  However, if $\alpha\in \k^\times\setminus\{1\}$, then the algebra is an Ore extension $\k[x_1,x_3,x_4][x_2;\varphi]$ where $\varphi(x_1)=(\alpha-1)x_1$, $\varphi(x_3)=x_3$ and $\varphi(x_4)=x_4$ and so, the algebra is regular of global dimension four. Moreover, the automorphism $\sigma|_L$ can be computed by factoring out the ideal $x_3R+x_4R$. 
				
\item	If $L$ is a line in $\P^3$ that meets $Q$ in one point counted with multiplicity two, then we choose coordinates $x_1, x_2, x_3, x_4$ such that $L=\V(x_1-x_2,x_3)$ and $Q=\V(x_1x_2)$.  By Proposition \ref{centralElts}, we may choose $x_1-x_2$ and $x_3$ to be central elements in $R$.  This gives seven possible defining relations: $x_3 x_i = x_i x_3$, for $i \in \{1,2,4\}$, and $(x_1-x_2)x_j = x_j(x_1-x_2)$, for $j\in \{1,2,3,4\}$. After eliminating some redundancies in the relations, we have the following four relations: 
\[\begin{matrix}
x_1x_2 = x_2x_1, & x_2x_3 = x_3x_2\\
x_1x_3 = x_3x_1, & x_3x_4 = x_4x_3
\end{matrix}\]
By \Cref{omega}, $\alpha x_1x_2 = x_1x_4-x_4x_1$. For the last relation, observe that $x_1x_4-x_4x_1=x_2x_4-x_4x_2$. This yields the final defining relation of $R$, $\alpha x_1x_2 = x_2x_4-x_4x_2$.  Mapping $x_1 \mapsto \sqrt{\alpha^{-1}}x_1$, $x_2 \mapsto \sqrt{\alpha^{-1}}x_2$, $x_3 \mapsto x_3$, $x_4 \mapsto \sqrt{\alpha}x_4$ gives the defining relations in (c). This algebra is the Ore extension $\k[x_1,x_2,x_3][x_4;\id,\delta]$ where $\delta(x_1)=-x_1x_2$, $\delta(x_2)=-x_1x_2$, $\delta(x_3)=0$; it may be verified that $\delta$ is a derivation and, hence, that $R$ is regular. We compute $\sigma|_L$ by factoring out the ideal $(x_1-x_2)R+x_3R$. 
				
\item	If $L$ is the line of intersection of the planes of $Q$, then there exists a choice of coordinates $x_1, x_2, x_3, x_4$ such that $L=\V(x_1,x_2)$ and $Q=\V(x_1x_2)$, and by \Cref{centralElts}, we may choose $x_1$ and $x_2$ to be central elements in $R$.   Thus, five of the six defining relations are determined.  It can be shown that $\alpha x_1 x_2 = x_3 x_4 - x_4 x_3 \in \k^\times \Omega$ for some $\alpha \in \k^\times$, which gives the final defining relation. Moreover, $x_1 \mapsto \sqrt{\alpha^{-1}}x_1$, $x_2 \mapsto \sqrt{\alpha^{-1}}x_2$, $x_3 \mapsto x_3$, $x_4 \mapsto x_4$ gives the desired defining relations.  This algebra is the Ore extension $\k[x_1,x_2,x_3][x_4;\id,\delta]$ where $\delta(x_1)=0$, $\delta(x_2)=0$, $\delta(x_3)=-x_1x_2$ and so, is regular.
				
To compute the action of the automorphism, we follow the methods in \cite[Proposition 2.6]{SV99} and set $Y_i'=x_i\otimes 1$ and $Z_i'=1\otimes x_i$ for all $i=1,...,4$ in the defining relations of the algebra.  This yields the (commutative) homogeneous coordinate ring $S$ corresponding to the point scheme $P$, $S(P)$, which is isomorphic to the commutative algebra $\k[Y_1',Y_2',Y_3',Y_4',Z_1',Z_2',Z_3',Z_4']$ with defining relations $$\begin{matrix} Y_1' Z_2' = Y_2' Z_1', & Y_2' Z_3' = Y_3' Z_2', \\ Y_1' Z_3' = Y_3' Z_1', & Y_2' Z_4' = Y_4' Z_2', \\ Y_1' Z_4' = Y_4' Z_1', & Y_3' Z_4' - Y_4' Z_3' = Y_1' Z_2'.  \end{matrix}$$
				
The subalgebra generated by the $Y_i'$ (respectively, the $Z_i'$) is the homogeneous coordinate ring of the left (respectively, right) projection of $P$ into $\P^3$ and is isomorphic to $S(P)$.  For each $S(P)[x_j^{-1}]$, set $Y_i=Y_i'/Y_j'$ and $Z_i=Z_i'/Z_j'$ for all $i \neq j$.  On each of these localizations, $\sigma$ is given by $\sigma(Y_i)=Z_i$.
				
For $S(P)[x_j^{-1}]$ where $j=1,2$, $\sigma$ is the identity.  For $S(P)[x_3^{-1}]$, we have $Y_2Z_1=Y_1Z_2$, $Y_4Z_1=Y_1Z_4$, $Y_4Z_2=Y_2Z_4$, $Z_1=Y_1$, $Z_2=Y_2$, $Z_4 = Y_4+Y_1Y_2$.  For $S(P)[x_4^{-1}]$, we have $Y_2Z_1=Y_1Z_2$, $Y_3Z_1=Y_1Z_3$, $Y_3Z_2=Y_2Z_3$, $Z_1=Y_1$, $Z_2=Y_2$, $Z_3=Y_3-Y_1Y_2$. The result follows.
				
\item	If $L$ is a line embedded in one of the planes of $Q$, then there exists a choice of coordinates $x_1, x_2, x_3, x_4$ such that $L=\V(x_1,x_3)$ and $Q=\V(x_1x_2)$.  By Proposition \ref{centralElts}, we may choose $x_1$ and $x_3$ to be central elements in $R$.   Thus, five of the six defining relations are determined and, it can be shown that $\alpha x_1 x_2 = x_2 x_4 - x_4 x_2$ for some $\alpha \in \k^\times$, which gives our final defining relation. Mapping $x_1 \mapsto \alpha^{-1}x_1$, $x_2 \mapsto x_2$, $x_3 \mapsto x_3$, $x_4 \mapsto x_4$ gives the desired defining relations.  This algebra is an Ore extension $\k[x_1,x_2,x_3][x_4;\id,\delta]$ where $\delta(x_1)=0$, $\delta(x_2)=-x_2x_1$, $\delta(x_3)=0$; it may be verified that $\delta$ is a derivation and, hence, that $R$ is regular. Moreover, retaining the same notation as above, the algebra $S(P)$ is isomorphic to the commutative algebra $\k[Y_1',Y_2',Y_3',Y_4',Z_1',Z_2',Z_3',Z_4']$ with defining relations $$\begin{matrix} Y_1' Z_2' = Y_2' Z_1', & Y_2' Z_3' = Y_3' Z_2', \\ Y_1' Z_3' = Y_3' Z_1', & Y_3' Z_4' = Y_4' Z_3', \\ Y_1' Z_4' = Y_4' Z_1', & Y_2' Z_4' - Y_4' Z_2' = Y_2' Z_1'.\end{matrix}$$				
				
For $S(P)[x_j^{-1}]$ where $j=1,3$, $\sigma$ is the identity.  For $S(P)[x_2^{-1}]$, we have $Y_3Z_1=Y_1Z_3$, $Y_4Z_1=Y_1Z_4$, $Y_4Z_3=Y_3Z_4$, $Z_1=Y_1$, $Z_3=Y_3$, $Z_4=Y_1+Y_4$.  For $S(P)[x_4^{-1}]$, we have $Y_2Z_1=Y_1Z_2$, $Y_3Z_1=Y_1Z_3$, $Y_3Z_2=Y_2Z_3$, $Z_1=Y_1$, $Z_3=Y_3$, $Z_2=Y_2-Y_2Y_1$.  The result follows.
\end{enumerate}	
\end{proof}	

\begin{cor}\cite[Corollary 2.13]{VVr}\label{tausigma}
The map $\tau \in \Aut(P)$ and it commutes with $\sigma$ on $P$.
\end{cor}
	
\begin{prop}\label{PLTwist}
The regular algebra $R$ is isomorphic to one of the following algebras if and only if the point scheme $P$ of $R$ is not the quadric $Q$.

\begin{enumerate}[(a)]
    
    \item   $R=\k[x_1,x_2,x_3,x_4]$ with defining relations $x_i^\tau x_j=x_j^\tau x_i$, for $1\leq i,j\leq 4$ for all $\tau \in \Aut(\P^3)$.  In this case, $R$ is a twist of the polynomial ring $S(\P^3)$ by $\tau$ so $(P,\sigma)=(\P^3,\tau)$.

    \item   $R=\k[x_1,x_2,x_3,x_4]$ with defining relations $$\begin{matrix} x_3^\tau x_1 = x_1^\tau x_3, & x_4^\tau x_2 = x_2^\tau x_4, \\ x_3^\tau x_2 = x_2^\tau x_3, & x_4^\tau x_1 = x_1^\tau x_4, \\ x_3^\tau x_4 = x_4^\tau x_3, & x_2^\tau x_1 = (\alpha - 1) x_1^\tau x_2,  \end{matrix}$$ where $\alpha \in \k^\times\setminus \{1,2\}$ and $\tau$ is given (with respect to a basis dual to $\{x_1, x_2, x_3, x_4\}$) by $$\tau \in \k^\times \begin{pmatrix} 1 & 0 & 0 & 0 \\ 0 & t_{22} & 0 & 0 \\ 0 & 0 & t_{33} & t_{34} \\ 0 & 0 & t_{43} & t_{44} \end{pmatrix},$$ where $t_{22} \in \k^\times$, $t_{ij}\in \k$ for $i,j\in \{3,4\}$, and $t_{33}t_{44}-t_{34}t_{43} \neq 0$.  In this case, $P=Q \cup L$ where $L=\V(x_3,x_4)$ (so $L$ intersects $Q$ at two distinct points), $\sigma|_Q=\tau$, and $\sigma|_L(x_1,x_2,0,0)=\tau\left( (\alpha-1)x_1,x_2,0,0\right)$.
                
    \item   $R=\k[x_1,x_2,x_3,x_4]$ with defining relations $$\begin{matrix} x_3^\tau x_1 = x_1^\tau x_3, & x_2^\tau x_1 = x_1^\tau x_2, \\ x_3^\tau x_2 = x_2^\tau x_3, & x_2^\tau x_4 - x_4^\tau x_2 = x_1^\tau x_2, \\ x_3^\tau x_4 = x_4^\tau x_3, & x_1^\tau x_4 - x_4^\tau x_1 = x_1^\tau x_2,  \end{matrix}$$ where $\tau$ is given (with respect to a basis dual to the $x_i$) by $$\tau \in \k^\times \begin{pmatrix} 1 & 0 & 0 & 0 \\ 0 & 1 & 0 & 0 \\ t_{31} & -t_{31} & t_{33} & 0 \\ t_{41} & t_{42} & t_{43} & 1 \end{pmatrix} \text{ or } \tau \in \k^\times \begin{pmatrix} 0 & 1 & 0 & 0 \\ 1 & 0 & 0 & 0 \\ t_{31} & -t_{31} & t_{33} & 0 \\ t_{41} & t_{42} & t_{43} & 1 \end{pmatrix},$$ where $t_{33} \in \k^\times$, $t_{ij}\in \k$.  In this case, $P=Q \cup L$ where $L=\V(x_1-x_2,x_3)$ (so $L$ intersects $Q$ at one point with multiplicity two), $\sigma|_Q=\tau$, and $\sigma|_L(x_1,x_1,0,x_4)=\tau(x_1,x_1,0,x_1+x_4)$.
    
    \item   $R=\k[x_1,x_2,x_3,x_4]$ with defining relations $$\begin{matrix} x_1^\tau x_2 = x_2^\tau x_1, & x_2^\tau x_3 = x_3^\tau x_2, \\ x_1^\tau x_3 = x_3^\tau x_1, & x_2^\tau x_4 = x_4^\tau x_2, \\ x_1^\tau x_4 = x_4^\tau x_1, & x_3^\tau x_4 - x_4^\tau x_3 = x_1^\tau x_2, \end{matrix}$$ where $\tau$ is given (with respect to a basis dual to the $x_i)$ by $$\tau \in \k^\times \begin{pmatrix} t_{11} & 0 & 0 & 0 \\ 0 & t_{22} & 0 & 0 \\ t_{31} & t_{32} & t_{33} & t_{34} \\ t_{41} & t_{42} & t_{43} & t_{44} \end{pmatrix} \text{ or } \tau \in \k^\times \begin{pmatrix} 0 & t_{12} & 0 & 0 \\ t_{21} & 0 & 0 & 0 \\ t_{31} & t_{32} & t_{33} & t_{34} \\ t_{41} & t_{42} & t_{43} & t_{44} \end{pmatrix} $$ where $t_{ij} \in \k$ and $t_{12}t_{21}+t_{11}t_{22}=t_{33} t_{44}-t_{34} t_{43} \neq 0$.  In this case, $P=Q\uplus L$ where $L=\V(x_1,x_2) \subset Q$ is the line of intersection the planes of $Q$, $\sigma|_Q=\tau$, and $\sigma \in \mathrm{Aut}(P)$ is uniquely determined by its action on the local rings $S(P)[x_i^{-1}]$, where $S(P)$ is the homogeneous coordinate ring of $P$ is as follows: on $S(P)[x_1^{-1}]$ and $S(P)[x_2^{-1}]$ we have that $\sigma = \tau^{-1}$, but on $S(P)[x_3^{-1}]$ we have $$\sigma(x_1,x_2,x_3,x_4)=\tau^{-1}\left(x_1,x_2,x_3,x_4-x_3^{-1}x_1x_2\right),$$ and on $S(P)[x_4^{-1}]$ we have $$\sigma(x_1,x_2,x_3,x_4)=\tau^{-1}\left( x_1,x_2,x_3+x_4^{-1}x_1x_2,x_4 \right).$$
                
    \item   $R=\k[x_1,x_2,x_3,x_4]$ with defining relations
	$$\begin{matrix} x_1^\tau x_2 = x_2^\tau x_1, & x_2^\tau x_3 = x_3^\tau x_2, \\ x_1^\tau x_3 = x_3^\tau x_1, & x_3^\tau x_4 = x_4^\tau x_3, \\ x_1^\tau x_4 = x_4^\tau x_1, & x_2^\tau x_4 - x_4^\tau x_2 = x_2^\tau x_1.  \end{matrix}$$ where $\tau$ is given (with respect to a basis dual to the $x_i)$ by $$\tau \in \k^\times \begin{pmatrix} 1 & 0 & 0 & 0 \\ 0 & t_{22} & 0 & 0 \\ t_{31} & 0 & t_{33} & 0 \\ t_{41} & t_{42} & t_{43} & 1 \end{pmatrix}$$ where $t_{ij}\in \k$ for $i\neq j$ and $t_{22},t_{33} \in \k^\times$. 
    In this case, $P=Q\uplus L$ where $L=\V(x_1,x_3) \subset \V(x_1) \subset Q$, $\sigma|_Q=\tau$, and $\sigma \in \mathrm{Aut}(P)$ is uniquely determined by its action on the local rings $S(P)[x_i^{-1}]$, where $S(P)$ is the homogeneous coordinate ring of $P$ is as follows: on $S(P)[x_1^{-1}]$ and $S(P)[x_3^{-1}]$ we have that $\sigma=\tau^{-1}$, but on $S(P)[x_2^{-1}]$ we have $$\sigma(x_1,x_2,x_3,x_4)=\tau^{-1}\left( x_1,x_2,x_3,x_1+x_4 \right),$$ and on $S(P)[x_4^{-1}]$ we have $$\sigma(x_1,x_2,x_3,x_4)=\tau^{-1}(x_1,x_2-x_4^{-1}x_2x_1,x_3,x_4).$$
    
\end{enumerate}
\end{prop}	
	
\begin{proof}
Using methods similar to \cite[Theorem 2.8]{SV99}, we classify the graded degree zero automorphisms of the algebras listed in Proposition \ref{PLAlg}.  Applying Corollary \ref{tausigma} to (a), (b) and (c) results in a matrix for $\tau$ that is determined by the hypotheses that $\tau\in \Aut(Q) \cap \Aut(L)$ and $\tau \mid_L \circ \ \sigma \mid_L = \sigma \mid_L \circ \ \tau \mid_L$ (since $\sigma|_Q=\id$).  The result follows by a computation, using Wolfram's Mathematica.
    
For (d) and (e), we apply $\tau$ to the defining relations of the respective algebras in Proposition \ref{PLAlg} rather than working with the commutativity of of $\tau$ and $\sigma$.
    
    In the case of (d), any element of $\Aut(Q) \cap \Aut(L)$ may be given (with respect to a basis dual to the $x_i$) by $$\tau \in \k^\times \begin{pmatrix} t_{11} & t_{12} & 0 & 0 \\ t_{21} & t_{22} & 0 & 0 \\ t_{31} & t_{32} & t_{33} & t_{34} \\ t_{41} & t_{42} & t_{43} & t_{44} \end{pmatrix},$$ where $t_{ij} \in \k$, $t_{11}t_{21}=t_{12}t_{22}=0$ and $(t_{12}t_{21}-t_{11}t_{22})(t_{34}t_{43}-t_{33}t_{44}) \neq 0$.  This matrix maps the span of the defining relations of the algebra back to itself if and only if $t_{12}t_{21}+t_{11}t_{22}+t_{34}t_{43}-t_{33}t_{44}=0$, so the result follows.

    In the case of (e), any element of $\Aut(Q) \cap \Aut(L)$ may be given (with respect to a basis dual to the $x_i$) by $$\tau \in \k^\times \begin{pmatrix} t_{11} & 0 & 0 & 0 \\ 0 & t_{22} & 0 & 0 \\ t_{31} & 0 & t_{33} & 0 \\ t_{41} & t_{42} & t_{43} & t_{44} \end{pmatrix},$$ where $t_{ij} \in \k$ for $i \neq j$ and $t_{ii} \in \k^\times$.  This matrix maps the span of the defining relations of the algebra back to itself if and only if $t_{11}t_{22} = t_{22} t_{44}$, so the result follows.
    
\end{proof}	
	
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
% END  SEC 1
% START  SEC 2
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%

\bigskip
%\newpage

\section{Regular Algebras whose Point Scheme is the Union of Two Planes}\label{sec3}

To motivate this section, we begin with an example from \cite[Lemma 5.1.1]{HTDiss} where the second author showed that there exist examples of AS-regular algebras of global dimension four such that their point schemes is exactly a rank-two quadric.

\begin{ex}\cite[Lemma 5.1.1]{HTDiss}\label{hung}
Consider the $\Bbbk$-algebra $A$ defined by $T(V)/W$ where $T(V)$ is the tensor algebra, $V$ is the vector space given by $V = \Bbbk x_1 \oplus \Bbbk x_2 \oplus \Bbbk x_3 \oplus \Bbbk x_4$ and $W$ is the ideal generated by the six relations given below.
\begin{align*}
x_{1} x_{2} - x_{2}x_{1}, \hspace{4mm}
& x_{2} x_{3} + x_{3}x_{2},\\
x_{1} x_{3} - x_{3}x_{1}, \hspace{4mm}
& x_{2} x_{4} + x_{4}x_{2}, \\
x_{1} x_{4} - x_{4}x_{1},\hspace{4mm}
&x_{3} x_{4} - x_{4}x_{3} +  x_{1} x_{2}.  \\
\end{align*}
\end{ex}

The point scheme of $A$ is exactly a rank-two quadric. In fact, Proposition \ref{prop1} below shows that when the point scheme of an algebra, satisfying the hypotheses of \Cref{HypsOnR}, is exactly a rank-two quadric, then a normal degree-one element always exists in $R$.

\begin{prop}\label{prop1}
If the point scheme $P$ of $R$ is $Q$, then there exists generators $x_1, \dots, x_4$ for $R$ such that $P = Q = \V(x_1x_2)$ and $R$ is isomorphic to one of the following algebras.

\begin{enumerate}[(a)]

\item   $R=\k[x_1,x_2,x_3,x_4]$ with defining relations

\begin{align*}
\alpha x_1^{\tau} x_2 = x_2^{\tau} x_1, \hspace{4mm}
& x_2^{\tau} x_3 = x_3^{\tau} x_2, \\ 
x_1^{\tau} x_3 = x_3^{\tau} x_1, \hspace{4mm}
&x_2^{\tau} x_4 = x_4^{\tau} x_2, \\ 
x_1^{\tau} x_4 = x_4^{\tau} x_1, \hspace{4mm}
& x_3^{\tau} x_4 - x_4^{\tau} x_3 = \beta x_1^{\tau} x_2,  \end{align*}

where $\alpha,\beta \in \k^\times$ and $\alpha\neq 1$.

\item   $R=\k[x_1,x_2,x_3,x_4]$ with defining relations 

\begin{align*}
x_1^\tau x_2 = x_2^\tau x_1, \hspace{4mm} & x_1^\tau x_3 = x_3^\tau x_1,\\  x_2^\tau x_3 - x_3^\tau x_2 = \alpha x_1^\tau x_2, \hspace{4mm} & x_1^\tau x_4 - x_4^\tau x_1 = \beta x_1^\tau x_2, \\x_2^\tau x_4 = x_4^\tau x_2, \hspace{4mm} & x_3^\tau x_4 = x_4^\tau x_3,
\end{align*}

where $\alpha, \beta \in \k^\times$.

\end{enumerate}

\begin{proof}

In this proof, we retain notation used in the proof of \Cref{OmegaNormal}.

\begin{enumerate}[(a)]

\item   By \Cref{omega}, we may choose $x_1, x_2, x_3$ such that $x_1^{\tau} x_3 = x_3^{\tau} x_1$ and $x_2^{\tau} x_3 = x_3^{\tau} x_2$ and moreover, we take $\Omega =\delta x_1^{\tau} x_2 $. Thus, we write $\alpha x_1^{\tau} x_2 = x_2^{\tau} x_1$ where $\alpha = 1-\alpha_{12}\delta$, $x_i^{\tau} x_4 - x_4^{\tau} x_i = \alpha_i x_1^{\tau} x_2$, and $i\in \{1,2,3\}$ with $\alpha_i \in \k$.  

If $\alpha = 1$, this algebra is isomorphic to the algebra $A=\k[y_1,y_2,y_3,y_4]$ with defining relations $y_1y_3=y_3y_1$, $y_2y_3=y_3y_2$, $y_1y_2=y_2y_1$ and $y_i y_4 - y_4 y_i = \alpha_i y_1 y_2$, where $i\in \{1,2,3\}$ and $\alpha_i \in \k$. The algebra $A$ is an Ore-extension of a polynomial ring on three variables given by the automorphism $\sigma(y_i)=0$, for all $i$, and the derivation $\gamma(y_1)=\gamma(y_2)=0$, $\gamma(y_3)=-\beta y_1y_2$ .  If, furthermore, $\alpha_i = 0$, for all $i$, then $A$ is the polynomial ring on four generators and $P=\P^3$.  If $\alpha_i \neq 0$ for some $i$, then $\alpha_i y_j - \alpha_j y_i$ are central, for $i,j \in \{1,2,3\}$ and so, by \Cref{quadline}, $Q \subset P$, which contradicts our hypothesis. Hence, we may assume $\alpha \neq 1$.  We rechoose $x_4$ such that $x_1^{\tau}x_4=x_4^{\tau}x_1$ and $x_2^{\tau}x_4 = x_4^{\tau} x_2$.  If, additionally, $x_3^{\tau} x_4 = x_4^{\tau} x_3$, then $R$ is a twist of the polynomial ring and its point scheme is not $Q$. Therefore, $x_3^{\tau}x_4 - x_4^{\tau} x_3 = \beta x_1^{\tau} x_2$, for $\beta \in \k^\times$.

\item If we choose $x_1,x_2,x_3,x_4$ such that $x_1^\tau x_2 = x_2^\tau x_1$, $x_1^\tau x_3 = x_3^\tau x_1$, $x_2^\tau x_3 - x_3^\tau x_2 = \alpha_{23} x_1^\tau x_2$, $x_2^\tau x_4 = x_4^\tau x_2$, and $x_3^\tau x_4 = x_4^\tau x_3$, where $\alpha_{23} \in \k^\times$, then the sixth relation is of the form $x_1 x_4 - x_4 x_1 = \alpha_{14} x_1x_2$, where $\alpha_{14} \in \k$. We write these relations as $Mx = 0$ where $M$ is a $6 \times 4$ matrix and $x^T = (x_1, x_2, x_3, x_3)$ and obtain the reduced scheme $P_{red}=Q$ where the ideal of $P_{red}$ is given by fifteen polynomials. Now, to determine the existence of points with multiplicity greater than one, by \cite[Definition \S I.5 \& Theorems 5.1 \& 5.3\footnote{While \cite[Theorem 5.1]{Hart} pertains to affine varieties, we know, using \cite[Corollary 2.3]{Hart}, that a projective variety can be covered by open sets which are then homeomorphic to affine varieties.}]{Hart}, we compute the $1\times 1$ minors of the Jacobian matrix of the fifteen polynomials of the ideal of $P_{red}$. This yields the possible points of higher multiplicity. 

If $\alpha_{14} \neq 0$, then the $1\times 1$ minors all vanish on $Q = \V(x_1,x_2)$ which implies that points of the form $(0, 0, a, b) \in \P^3$ might be of higher multiplicity. Next, we apply Bertini's Theorem \cite[Theorem 8.18]{Hart}, and intersect $Q$ with a complementary-dimensional scheme, that is, a 1-dimensional scheme, which is given by the line $L_n=\V(x_1-x_2,n x_3 - x_4)$, where $n\in \P^1$. The point of intersection of $L_n$ and $\V(x_1,x_2)$ is $(0,0,1,n)$, when $n \in \k$, or $(0,0,0,1)$, when $n$ is a point at infinity. We determine the vector space dimension of the localization of the local ring $\mathcal{O}_{Q \cap L_n}$ at each of these points to compute their multiplicities \cite{Hart}. We obtain a ring isomorphic to $\k[x]/\langle x^2 \rangle$ and so, the point has multiplicity two. This implies that the points on $\V(x_1,x_2)$ are multiple points only as a consequence of being in the intersection of two irreducible components of $P$, $\V(x_1)$ and $\V(x_2)$, explicitly. Taking $\alpha=\alpha_{23}$ and $\beta=\alpha_{14}$ gives the defining relations as described.  This algebra is an extension of the polynomial ring $\k[x_2,x_3,x_4]$ by the normal element $x_1$ and is thus, regular.

Otherwise, $\alpha_{14}=0$, and we show that, in this case, we have a contradiction. Observe that the $1 \times 1$ minors of the corresponding Jacobian matrix vanish on $\V(x_1,x_2)$ and on $\V(x_1,x_4)$.  Using $L_n$ to examine the multiplicity of the points $\V(x_1,x_2)$ as before, we find that all points on $\V(x_1,x_2)$ have multiplicity two (as a consequence of the intersection of the irreducible components of $P$) except the point $(0,0,1,0)$, which has multiplicity three. To further analyze the point $(0, 0, 1, 0)$, consider the points on $\V(x_1,x_4)\setminus\V(x_1,x_2)$ that are of the form $(0,1,n,0)$, where $n\in \k$.  In a similar manner as above, we examine the multiplicity of these points using the line $K_n=\V(x_3- n x_2,x_1-x_4)$.  In doing so, we find that the coordinate ring of $P \cap K_n$ is isomorphic to $\k[x]/\langle x^2 \rangle$ and so, the points have multiplicity two. This implies that the point $(0, 0, 1, 0)$ which lies on the line $\V(x_1,x_4)$ is a point on an embedded line in $Q$. Thus, the point scheme of $R$ is $Q \uplus \V(x_1,x_4)$, which contradicts our hypothesis that $P = Q$.

\end{enumerate}

\end{proof}
\end{prop}

\begin{rmk}
If $P = Q$ and there is a normal element $z \in R_1$, then $z$ is also regular as $R$ is a domain. Moreover, by \cite[Theorem 8.16(ii)]{ATV2}, the normal element $z^{\tau} \in R_1$ for an automorphism $\tau$ is central in $R^{\tau}$. As a result, in this case, $R$ is a central extension of an AS-regular of global dimension three. These central extensions are well-understood algebras (see e.g., \cite{lBSV}), and by Proposition \ref{prop1}, the algebra $R$ will always be such an algebra.
\end{rmk}

\begin{ex}
The map $\tau$ in Proposition \ref{prop1}(a) is given by the automorphism $$\tau \in \k^\times \begin{pmatrix} t_{11} & 0 & 0 & 0 \\ 0 & t_{22} & 0 & 0 \\ 0 & 0 & t_{33} & t_{34} \\ 0 & 0 & t_{43} & t_{44} \end{pmatrix},$$ where $t_{ij} \in \k$ and $t_{11}t_{22}=t_{33}t_{44}-t_{34}t_{43} \neq 0$.  If $\alpha = -1$, then $\tau$ may also be given by $$\tau \in \k^\times \begin{pmatrix} 0 & t_{12} & 0 & 0 \\ t_{21} & 0 & 0 & 0 \\ 0 & 0 & t_{33} & t_{34} \\ 0 & 0 & t_{43} & t_{44} \end{pmatrix},$$ where $t_{ij} \in \k$ and $t_{12}t_{21}=t_{34}t_{43}-t_{33}t_{44} \neq 0$.

If $\tau \in \Aut(Q)$, then $\tau$ is given by $$\begin{pmatrix} t_{11} & 0 & 0 & 0 \\ 0 & t_{22} & 0 & 0 \\ t_{31} & t_{32} & t_{33} & t_{34} \\ t_{41} & t_{42} & t_{43} & t_{44} \end{pmatrix} \text{ or }\begin{pmatrix} 0 & t_{12} & 0 & 0 \\ t_{21} & 0 & 0 & 0 \\ t_{31} & t_{32} & t_{33} & t_{34} \\ t_{41} & t_{42} & t_{43} & t_{44}\end{pmatrix}.$$  The first matrix maps the span of the defining relations of $R$ back to itself if and only if $t_{31}=t_{32}=t_{41}=t_{42}=0$ and $t_{11}t_{22}+t_{34}t_{43}-t_{33}t_{44}=0$. The second matrix maps the span of the defining relations of $R$ back to itself if and only if $(\alpha^2-1)t_{12}t_{21}=0$, $t_{31}=t_{32}=t_{41}=t_{42}=0$ and $t_{12}t_{21}-t_{34}t_{43}+t_{33}t_{44}=0$.  Since $\alpha \neq 1$ and $t_{12}t_{21} \neq 0$ it follows that $\alpha=-1$.
\end{ex}

\begin{rmk}
The 2-parameter family of algebras obtained in Proposition \ref{prop1} does not explicitly have the algebra $A$ of Example \ref{hung} as a member. However, $A$ can be obtained from the 2-parameter family via the twisting system defined in \cite[Definition 2.1]{HV} given by $t_n(a_1 \dots a_m) = \Pi_{j=1}^m \tau^{1-j-n}t^n\tau^{j-1}(a_j)$ where $m \in \mathbb{N}$ and $a_j \in T(V)_1$. Take 
\[t=\begin{pmatrix} a & 0 & 0 & 0 \\ 0 & 1 & 0 & 0 \\ 0 & 0 & 1 & 0 \\ 0 & 0 & 0 & 1 \end{pmatrix}, \mbox{ and } \hspace{1mm}
\tau^{-1}=\begin{pmatrix} -q^{-1} & 0 & 0 & 0 \\ 0 & -q & 0 & 0 \\ 0 & 0 & -1 & 0 \\ 0 & 0 & 0 & 1 \end{pmatrix},\] to obtain the algebra $A$.
\end{rmk}

\bibliography{BiblioTHCR}
\bibliographystyle{plain}

\end{document}